When separated by distance (d), identically charged objects, A and B exert a force of magnitude F on each other. if you reduce the charge of A to one half its original value, and the charge of B to one fourth, and reduce the distance between objects by half, what will be the new force that they will exert on each other.

Choices
a. 2F
b. F/2
c. 4F
d. F/4
e. F

Please explain your answer with steps since it's a bit confusing.

Thanks
Answer #1

Option B is correct

Intially let F be= (k*q1*q2)/(r^2)

New q1= (1/2)*q1

New q2= (1/4)*q2

New distance= (1/2)

Now new force will be F'= [k*(q1/2)*(q2/4)]/[(r/2)^2]

F'= [(1/2)*k*q1*q2]/(r^2)

New force= (1/2)*F